1. find the domains of f(x)= √x+7 / x+4= g(x) = 3 / |x| + 2 
  2. Draw the graph of f(x) =x3. then use it together with some transformation to sketch the graph of g(x) =(x - 3)3 - 4.
1
Expert's answer
2021-10-06T17:59:08-0400

1. Let us find the domains of the functions  "f(x)=\frac{ \sqrt{x+7}}{ x+4}" and "g(x) = \frac{3}{|x| + 2}." The domain of function "f" consists of all real numbers "x" such that "x+7\ge0" and "x\ne-4." Therefore, this domain is equal to "[-7,-4)\cup(-4,+\infty)." Since "|x|+2>0" for each "x\in\R," we conclude that the domain of the function "g" is the set "\R" of all real numbers.


2. Let us draw the graph of "f(x) =x^3:"



Then use it together with the shift 3 units to the right and 4 units downward to sketch the graph of the function "g(x) =(x - 3)^3 - 4:"